OpenStudy (anonymous):

completing the square x^2-4x+1=0

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

x^2-4x+1=0 x^2-4x+4-3=0 (x^2-4x+4)=3 (x-2)^2=3 (x-2)=+- sqrt{3} x=2+-sqrt(3)